OpenStudy (anonymous):

If the given polynomial has a special factorization, factor the polynomial: 25x^2-16

OpenStudy (anonymous):

difference of two squares

OpenStudy (saifoo.khan):

(5x + 4) (5x - 4)

OpenStudy (anonymous):

\[a^2-b^2=(a+b)(a-b)\] in this case \[a=5x,b=4\] gives \[25x^2-16=(5x+4)(5x-4)\]

OpenStudy (anonymous):

difference of two squares